# Breakdown of broken-symmetry approach to exchange interaction

### v1

@@ II.2 Exact vs BS calculations of $J$
where $\bm{S}$ is the classical spin vector, and the last expression assumes a small $\theta$. Within the quantum mechanical treatment, the rotation of the quantum spins of the BS-HF state is achieved by $\hat{R}(\theta)|\Psi_{\text{UHF}}\rangle$ with
@@
By comparing the classical and quantum mechanical energies, we obtain $J$. See for the detailed expressions of the energy Appendix B.
Figure 3(a) shows the calculated exchange parameter in the function of $t^{\prime}/U$. The exact $J$ obtained from the full CI calculations (solid), the broken-symmetry $J_{\text{BS}}$ (dashed), and the magnetic force theorem $J_{\text{MFT}}$ (symbols) exhibit similar behavior, whereas the approximate $J$’s are quantitatively and, under certain ranges of parameters, qualitatively different from the exact one. $J_{\text{BS}}$ tends to overestimate the ferromagnetic contribution due to the contamination of the ferromagnetic component of about 40-50 % in the UHF energy (see Appendix A.2.2). The spin contamination in the BS-HF wave function makes the variational parameters (molecular orbital coefficients) not fully optimal for the description of the ground antiferromagnetic state 444 One may think that the breakdown of the BS approach may be cured by using a spin symmetry adapted state instead of the UHF state (see Appendix A.2.3). We confirm with our toy model that such an approach is far from satisfactory in reproducing the exact $J$. To improve the BS approach, the use of a multiconfigurational approach is decisive. In practice, similar attempts have been made using DFT, while clear improvement of $J$ of real molecules has not been seen [86]. . Moreover, the number of the variational parameters for the BS-HF wave function is less than that for the exact solution, leading to a poor description of the antiferromagnetic state with the BS-HF wave function. For some range of $t^{\prime}/U$, one can also see that $J_{\text{BS}}$ qualitatively differs (opposite sign) from the exact $J$: Within $0.03\lesssim t^{\prime}/U\lesssim 0.05$ and $0.22\lesssim t^{\prime}/U\lesssim 0.26$ for $U^{\prime}/U=1$ ($0.21\lesssim t^{\prime}/U\lesssim 0.24$ for $U^{\prime}/U=0$), the exact $J$ becomes antiferromagnetic, whereas $J_{\text{BS}}$ is ferromagnetic. The discrepancy is enlarged with the increase of the Coulomb repulsion on the bridging site, $U^{\prime}$, implying that the mean-field description is not adequate. The static electron correlation effect is crucial to derive accurate $J$.
We can see the importance of the electron correlation for the description of $J$ by modulating the covalency via $\Delta$. From a detailed analysis of the present model, we have demonstrated that the static electron correlation effect is enhanced for small $\Delta$ [33]. Figures 3(b), (c) show that $J_{\text{BS}}$ deviates from exact $J$ when $\Delta$ lowers and the covalency effect enhances: The discrepancy is further enhanced by turning on $U^{\prime}$.
@@
$J_{\text{MFT}}$ shows opposite tendency to $J_{\text{BS}}$. First, $J_{\text{MFT}}$ overestimates the antiferromagnetic contribution [Fig. 3(a)]: $J_{\text{MFT}}$ is ferromagnetic in a smaller range of $t^{\prime}/U$ than the exact $J$ does and exhibits larger antiferromagnetic contribution. Second, the description of $J_{\text{MFT}}$ becomes poor when the covalency effect becomes stronger by reducing $U^{\prime}$ or increasing $t$’s [Figs. 3(b)-(d)]. Since the magnetic force theorem calculations assume well-localized classical spins, this method would not work well when the covalency effects are strong.
@@
The present analysis shows that the BS approache and the MFT one fail when the magnetic electrons are delocalized over the bridging ligand, realized at sufficiently small $\Delta$. Small values of $\Delta$ are not unusual when the metal and ligand orbitals are close to resonance, e.g., in iron-sulfur complexes [45]. Also, as emphasized above, finite and even large $t^{\prime}$ is not surprising in real materials.
@@
The breakdown of the BS approach for the calculation of $J$ ultimately originates from the lack of sufficient flexibility of the BS-HF determinant. Indeed, while CI and HF wave functions for ferromagnetic states coincide in our model, they differ for the antiferromagnetic states through fewer variational parameters in the BS-HF treatment (see Appendix A.2.2). The absence of approximation in the derivation of exact states excludes any doubts that the discrepancy shown by the BS approach is only due to the intrinsic drawback of the latter. In particular, the significance of adequate treatment of electron correlation in the antiferromagnetic states is revealed. To eliminate this intrinsic drawback, the BS approach should be extended as suggested below.
@@
III Extension of the BS approach for the exchange parameters
A reliable multiconfigurational calculation of exchange parameters based on Hartree-Fock orbitals as e.g., within the CASSCF/CASPT2 approximation [46], poses difficulties, especially, for large molecules. The main drawback of this approach is an insufficient account of dynamical correlation resulting in underestimation of the intersite electron transfer and overestimation of on-site electron repulsion. On the contrary, the dynamical correlation is intrinsically contained in DFT, CC and GW, which is the reason why the BS approaches became a standard tool for the calculation of $J$. Here we suggest an extension of this approach in order to cure its intrinsic drawbacks described above. To this end, we propose a minimal multiconfigurational version involving electronic configurations built on several active orbitals derived from a preliminary BS calculation. To exemplify this methodology, we consider the simplest case of two unpaired electrons in equivalent magnetic (metal) sites, the extension to other situations being straightforward.
III.1 Spin polarization contribution
@@
This equation is a particular case of the Yamaguchi’s formula (3) when the spin contamination comes from a single pair of non-orthogonal magnetic orbitals [47]. The total exchange parameter consists of the contributions described in the previous section ($J$) and the spin-polarization contribution (7):
@@
III.2 The active orbitals
To overcome the drawbacks of the BS approach, we should be in line with the results of the previous section, i.e. identify a set of effective magnetic and bridging ligand orbitals (active orbitals) and undertake a multiconfigurational calculation on their basis. Then from the energies of the lowest $S=0$ and $S=1$ states the parameter $J$ is extracted. Following the previous section, these active orbitals are chosen in a form which allows the expansion of the BS magnetic orbitals $\psi^{\sigma}$ solely in their basis. The construction of these active orbitals is shown below for a simplest system of two equivalent magnetic (metal) sites $i$=1,2 with one unpaired electron, $S_{i}$=1/2. Furthermore the complex is supposed to possess a mirror symmetry under which the BS orbitals $\psi^{\sigma}$ pass into each other, $\sigma_{\text{h}}\psi^{\uparrow}=\psi^{\downarrow}$ [Fig. 4(a)], whereas the restricted molecular orbitals $\psi_{i},\psi_{\mu}$} are either even or odd with respect to this transformation, i.e., are characterized by indices $g$ and $u$ respectively  [^5] In the case when the complex possesses a two-fold rotational symmetry interchanging the magnetic centres, the indices $g$ and $u$ are replaced by the irreducible representations $a_{1}$ and $a_{2}$ of the C2 symmetry group, respectively. .
@@ III.2 The active orbitals
whereas from the normality of $\psi^{\sigma}$ and orthogonality of $\phi_{M_{1}}$ and $\phi_{M_{2}}$ we extract the coefficients $c_{1}$ and $c_{2}$, Eq. (43), entering the decomposition (10). As Eqs. (14) and (43) show, all coefficients in this decomposition are expressed via $d^{\text{o}}_{g}$ and $d^{\text{e}}_{u}$, which together with the coefficients {$a_{\mu g}$} and {$a_{iu}$} are the variational parameters defining the two active ligand orbitals in Eq. (11). Once these are found from an optimization procedure specified below, all expansion coefficients in Eqs. (10) can be calculated, which allows via the knowledge of $\psi^{\sigma}$ to determine straightforwardly from these equations for the active magnetic orbitals $\phi_{M_{1}}$ and $\phi_{M_{2}}$ [Eq. (44)].
In the case of non-equivalent magnetic centers, the two BS magnetic orbitals $\psi^{\sigma}$ will not be related by symmetry anymore. Therefore, there is no reason to expect that they will decompose through common active ligand orbitals as in Eq. (10) but rather through different ones. Passing from two different active ligand orbitals of each type to their orthogonal combinations, we can decompose $\psi^{\uparrow}$ and $\psi^{\downarrow}$ into six common active orbitals: two magnetic, two ligand of doubly occupied type and two ligand of empty type. This scheme is straightforwardly generalized to several unpaired electrons on magnetic sites and more than two magnetic centers. In the absence of symmetry we will have to define for each magnetic electron ($n$) one magnetic ($\phi_{M_{n}}$) and two ligand ($\phi_{ln}^{\text{o}}$ and $\phi_{ln}^{\text{e}}$) active orbitals. The latter are written in analogy to Eq. (11) as follows:
@@
Besides serving as a tool to calculate $J$, this methodology allows to extract the magnetic and effective ligand orbitals in a strict variational way. The existent approaches merely identify them with some Wannier orbitals constructed from arbitrarily chosen group of molecular/band orbitals [49], an a priori unjustifiable and often unreliable procedure, especially, for large ligands.
III.3 A minimal version of MC extension
@@
where $N_{\text{d}}$ is the number of doubly occupied restricted orbitals in the DFT calculation. Since all configurations $\Phi_{I}$ in Eq. (18) involve either the l.h.s. determinant as is or with few electrons removed from a relatively small number of orbitals $\psi_{ln}^{\text{o}}$, $\bar{\psi}_{ln}^{\text{o}}$, we can describe them via few holes added to the core determinant (19). Since in addition we are not re-optimizing the orbitals $\psi_{\mu}$ during the MC calculation, the determinant in the r.h.s. represents a “vacuum function” for the added holes. At the same time the occupation of BS orbitals (their combinations $\psi_{m}^{\text{BS}}$) and empty restricted orbitals $\psi_{i}$ is described in the electronic representation. In this way the wave functions $\Phi_{I}$ in the expansion (18) involve explicitly only few electrons in the orthogonal BS orbitals and empty restricted orbitals and few holes in the doubly occupied restricted orbitals. Details of such electron-hole description are given in Appendix C.2.
@@
In this work we do not discuss the implementation of the proposed approach. One should mention, however, that the feasibility of this scheme depends on the evaluation of off-diagonal matrix elements in Eq. (20). It is a quite straightfoward procedure for HF, CC and even GW approaches. As for DFT, the evaluation of $H_{IJ}$ can only be done indirectly, within an uncontrolled approximation. Given the popularity of DFT calculations, we review below the MC DFT approaches used to date.
III.3.1 MC DFT approaches
Several approaches to the MC DFT have been developed in the past. These include re-definition of Kohn–Sham (KS) theory to include multiconfigurational reference wave function from the start [52, 53], a range-separation of the electron–electron interaction into a short-range part described by a local correlation potential and a long-range part described by the correlation arising from the MC expansion [54, 55, 56], a method based on a local scaling factor of the DFT correlation energy [57] and more complicated methods for the balanced treatment of MC and DFT correlation effects [58], as well as methods based on a correlation separation using the LDA correlation energy density [59, 60, 61]. As a less rigorous approach, a reparametrization of the XC functional in the context of an MC expansion [62, 63] and the rescaling of the matrix elements of the CI matrix constructed in the presence of the XC potential by empirical coefficients [64] have been proposed. Another actively developed approach is the so-called ensemble-referenced Kohn-Sham (REKS) method [65, 66, 67], in which the variational entity is an ensemble density expanded as a linear combination of densities corresponding to individual determinants. We note that all these approaches can be applied for our problem in a slightly modified form.
Concerning Eq. (20), within DFT the matrix elements are supposed to be rescaled by empirical coefficients [64]. This approach is closely related to the actively used nowadays ROCIS method, a single-configuration multireference approach for evaluation of electronic excitation of inner shells [68, 69], which also employs the rescaling of CI matrix elements using empirical factors. As starting point in the self-consistent calculation, we identify $\phi_{M_{n}}$ with $\psi_{n}$ which are maximally close to the original $\psi_{m}^{\text{BS}}$, obtained from the latter via e.g. a Löwdin orthogonalization ($c_{nm}=\delta_{nm}$, $c_{nm}^{\text{o,e}}=0$); the orbitals $\phi_{ln}^{\text{o,e}}$ are identified with corresponding restricted orbitals having maximal weight in the pairs of neighbor (overlapping) $\psi_{m}^{\text{BS}}$s.

### v2

@@ IV.2 Magnetic force theorem
where $\bm{S}$ is the classical spin vector, and the last equality assumes a small $\theta$. Within the quantum mechanical treatment, the rotation of the quantum spins of the BS-HF state is achieved by $\hat{R}(\theta)|\Psi_{\text{BS}}\rangle$ with
@@
By comparing the classical and quantum mechanical energies, we obtain $J_{\text{MFT}}$. The expressions for the energy and $J_{\text{MFT}}$ are given in Appendix B.
@@
Fig. 2 (symbols) shows the resulting exchange parameter as function of $t^{\prime}/U$ and $\Delta/U$. One can see that $J_{\text{MFT}}$ shows opposite tendencies to $J_{\text{BS}}$ and $J_{\text{SA}}$. First, $J_{\text{MFT}}$ overestimates the antiferromagnetic contribution [Fig. 2(a)]. Thus MFT exchange parameter is ferromagnetic in a smaller range of $t^{\prime}/U$ than the exact $J$ and exhibits larger antiferromagnetic contribution than the latter. Second, the description of $J_{\text{MFT}}$ becomes poor when the covalency effects becomes stronger by reducing $U^{\prime}$ or increasing $t$’s [Figs. 2(b)-(d)]. Since the magnetic force theorem calculations assume well-localized classical spins, this method would not work well when the covalency effects are strong.
IV.3 Spin-flip approach
Another single-reference method that allows to maintain the symmetry of the involved spin states is the spin-flip time-dependent DFT approach (SF-TDDFT) [50]. It starts from a highest-spin configuration within the subsystem of unpaired electrons ($S_{\text{max}}$) and considers a set of one-electron excitations accompanied by the reversal of one electron spin, which after configuration mixing results in a number excited terms with spin $S_{\text{max}}-1$. The closest of them to the reference one can be used for the extraction of exchange parameters. For instance, in the case of two magnetic centers ($A$ and $B$), we have
$J_{AB}\;S_{\text{max}}=E\big{(}S_{\text{max}}\big{)}-E\big{(}S_{\text{max}}-1\big{)}.$ (10)
This approach can be applied straightforwardly also to the cases when the state with $S_{\text{max}}$ is not the ground one, which makes it very useful for the calculation of exchange parameters in exchange coupled complexes [51, 52, 53]. It can be equally applied to complexes with a large number of magnetic sites, for which it is especially efficient even in comparison with BS DFT approaches [54].
One clear advantage of the SF TDDFT method is the correct description of single-electron excitation energies, which implicitly contain the screening of electronic interaction due to dynamical correlation, first of all, the strong reduction of the $U$ parameter in (1) compared to its bare HF value. Failure to account properly for the screening of $U$ is the major reason for the strong underestimation of antiferromagnetic contribution to $J$ by CASSCF methods. However, the SF TDDFT has an intrinsic drawback of not taking into account many single-electron spin-flip excitations, which may contribute to $J$ Figure 5 show the spin-flip excitation from the reference $S=1$ configuration which might be included in a SF TDDFT calculation when applied to the two-site model (1). We see that the requirement of a proper account of spin symmetry of excited singlets rules out completely the excitation involving ligand orbitals [in Fig. 5 a hole picture of the ground and SF configurations is employed for simplicity]. This will certainly affect the value of $J$ when the ligand-metal covalency is not small.
Figure 5: Electron configurations within the spin-flip approach. (a) Reference high-spin state. (b) Spin-flip states.
IV.4 Missing contributions to $J$ in the single-reference approaches
@@
The present analysis shows that the single-reference based approaches fail to provide reliable $J$ values when the magnetic electrons are significantly delocalized over the bridging ligand, a situation realized at sufficiently small $\Delta$. Small values of $\Delta$ arise when the metal and ligand orbitals are close to resonance as, e.g., in iron-sulfur complexes [55, 33]. Weaker but non-negligible delocalization occurs in other cases, such as copper-oxygen systems [33] and iron-oxygen complexes [56]. Also, as emphasized in Sec. II, finite and even large $t^{\prime}$ is not surprising in real materials [33].
@@
The breakdown of the single-reference approaches for the calculation of $J$ clearly originates from the lack of sufficient flexibility of the trial wave functions. Indeed, while CI and HF wave functions for ferromagnetic states coincide in our model, they differ for the antiferromagnetic ($S=0$) states through fewer variational parameters in the latter treatment (see Appendix A.2.2). The absence of approximations in the derivation of exact states rules out any doubts that the discrepancy shown by the BS approach is due to its intrinsic drawback. This consists ultimately in the lack of adequate treatment of electron correlation involving configurations of bridging ligand type. To eliminate this drawback, the BS approach should be extended as suggested below.
@@
V Extension of the BS approach for the exchange parameters
A reliable multiconfigurational calculation of exchange parameters based on Hartree-Fock orbitals, e.g., within the CASSCF/CASPT2 approximation [57], poses difficulties, especially for large molecules. The main drawback of this approach is an insufficient account of dynamical correlation, resulting in an underestimation of the intersite electron transfer and an overestimation of on-site electron repulsion. On the contrary, the dynamical correlation is intrinsically contained in DFT, CC and GW, which is the reason why the BS approaches became a standard tool for the calculation of $J$. Here we suggest an extension of this approach in order to cure its intrinsic drawbacks described above. To this end, we propose a minimal multiconfigurational version involving electronic configurations built on several active orbitals derived from a preliminary BS calculation. To exemplify this methodology, we consider the simplest case of two unpaired electrons in equivalent magnetic (metal) sites, the extension to other situations is straightforward.
V.1 Spin polarization contribution
@@
This equation is a particular case of Yamaguchi’s formula (4) when the spin contamination comes from a single pair of non-orthogonal magnetic orbitals [58]. The total exchange parameter consists of the contributions described in the previous section ($J$) and the spin-polarization contribution (12):
@@
V.2 The active orbitals
To overcome the drawbacks of the BS approach, we should be in line with the results of the previous section, i.e. identify a set of effective magnetic and bridging ligand orbitals (active orbitals) and undertake a multiconfigurational calculation on their basis. The n from the energies of the lowest $S=0$ and $S=1$ states the parameter $J$ is extracted. Following the previous section, these active orbitals are chosen in a form that allows the expansion of the BS magnetic orbitals $\psi^{\sigma}$ solely in their basis. The construction of these active orbitals is shown below for the simplest system of two equivalent magnetic (metal) sites $i$=1,2 with one unpaired electron, $S_{i}$=1/2. Furthermore, the complex is supposed to possess a mirror symmetry under which the BS orbitals $\psi^{\sigma}$ pass into each other, $\sigma_{\text{h}}\psi^{\uparrow}=\psi^{\downarrow}$ [Fig. 6(a)], whereas the restricted molecular orbitals $\psi_{i},\psi_{\mu}$} are either even or odd with respect to this transformation, i.e., are characterized by indices $g$ and $u$ respectively  [^6] In the case when the complex possesses a two-fold rotational symmetry interchanging the magnetic centers, the indices $g$ and $u$ are replaced by the irreducible representations $a_{1}$ and $a_{2}$ of the C2 symmetry group, respectively. .
@@ V.2 The active orbitals
whereas from the normality of $\psi^{\sigma}$ and orthogonality of $\phi_{M_{1}}$ and $\phi_{M_{2}}$ we extract the coefficients $c_{1}$ and $c_{2}$, Eq. (50), entering the decomposition (15). As Eqs. (19) and (50) show that all coefficients in this decomposition are expressed via $d^{\text{o}}_{g}$ and $d^{\text{e}}_{u}$, which together with the coefficients {$a_{\mu g}$} and {$a_{iu}$} are the variational parameters defining the two active ligand orbitals in Eq. (16). Once these are found from an optimization procedure specified below, all expansion coefficients in Eqs. (15) can be calculated, which allows via the knowledge of $\psi^{\sigma}$ to determine straightforwardly from these equations for the active magnetic orbitals $\phi_{M_{1}}$ and $\phi_{M_{2}}$ [Eq. (51)].
In the case of non-equivalent magnetic centers, the two BS magnetic orbitals $\psi^{\sigma}$ will not be related by symmetry anymore. Therefore, there is no reason to expect that they will decompose through common active ligand orbitals as in Eq. (15) but rather through different ones. Passing from two different active ligand orbitals of each type to their orthogonal combinations, we can decompose $\psi^{\uparrow}$ and $\psi^{\downarrow}$ into six common active orbitals: two magnetic, two ligands of doubly occupied type and two ligands of empty type. This scheme is straightforwardly generalized to several unpaired electrons on magnetic sites and more than two magnetic centers. In the absence of symmetry, we will have to define for each magnetic electron ($n$) one magnetic ($\phi_{M_{n}}$) and two ligands ($\phi_{ln}^{\text{o}}$ and $\phi_{ln}^{\text{e}}$) active orbitals. The latter are written in analogy to Eq. (16) as follows:
@@
Besides serving as a tool to calculate $J$, this methodology allows to extract the magnetic and effective ligand orbitals in a strict variational way. The existent approaches merely identify them with some Wannier orbitals constructed from arbitrarily chosen group of molecular/band orbitals [60], an a priori unjustifiable and often unreliable procedure, especially for large ligands.
V.3 A minimal version of MC extension
@@
where $N_{\text{d}}$ is the number of doubly occupied restricted orbitals in the DFT calculation. Since all configurations $\Phi_{I}$ in Eq. (23) involve either the l.h.s. determinant as is or with few electrons removed from a relatively small number of orbitals $\psi_{ln}^{\text{o}}$, $\bar{\psi}_{ln}^{\text{o}}$, we can describe them via few holes added to the core determinant (24). Since, in addition, we are not re-optimizing the orbitals $\psi_{\mu}$ during the MC calculation, the determinant in the r.h.s. represents a “vacuum function” for the added holes. At the same time, the occupation of BS orbitals (their combinations $\psi_{m}^{\text{BS}}$) and empty restricted orbitals $\psi_{i}$ are described in the electronic representation. In this way, the wave functions $\Phi_{I}$ in the expansion (23) involve explicitly only a few electrons in the orthogonal BS orbitals and empty restricted orbitals, and a few holes in the doubly occupied restricted orbitals. Details of such electron-hole description are given in Appendix C.2.
@@
In this work, we do not discuss the implementation of the proposed approach. One should mention, however, that the feasibility of this scheme depends on the evaluation of off-diagonal matrix elements in Eq. (25). It is a quite straightforward procedure for HF, CC, and even GW approaches. As for DFT, the evaluation of $H_{IJ}$ can only be done indirectly, within an uncontrolled approximation. Given the popularity of DFT calculations, we review the MC DFT approaches used to date below.
V.3.1 MC DFT approaches
Several approaches to the MC DFT have been developed in the past. These include a re-definition of Kohn–Sham (KS) theory to include multiconfigurational reference wave function from the start [63, 64], a range-separation of the electron-electron interaction into a short-range part described by a local correlation potential and a long-range part described by the correlation arising from the MC expansion [65, 66, 67], a method based on a local scaling factor of the DFT correlation energy [68] and more complicated methods for the balanced treatment of MC and DFT correlation effects [69], as well as methods based on a correlation separation using the LDA correlation energy density [70, 71, 72]. As a less rigorous approach, a reparametrization of the XC functional in the context of an MC expansion [73, 74], and the rescaling of the matrix elements of the CI matrix constructed in the presence of the XC potential by empirical coefficients [75] have been proposed. Another actively developed approach is the so-called ensemble-referenced Kohn-Sham (REKS) method [76, 77, 78], in which the variational entity is an ensemble density expanded as a linear combination of densities corresponding to individual determinants. We note that all these approaches can be applied to our problem in a slightly modified form.
Concerning Eq. (25), within DFT, the matrix elements are supposed to be rescaled by empirical coefficients [75]. This approach is closely related to the actively used nowadays ROCIS method, a single-configuration multireference approach for evaluation of electronic excitation of inner shells [79, 80], which also employs the rescaling of CI matrix elements using empirical factors. As starting point in the self-consistent calculation, we identify $\phi_{M_{n}}$ with $\psi_{n}$ which are maximally close to the original $\psi_{m}^{\text{BS}}$, obtained from the latter via e.g. a Löwdin orthogonalization ($c_{nm}=\delta_{nm}$, $c_{nm}^{\text{o,e}}=0$); the orbitals $\phi_{ln}^{\text{o,e}}$ are identified with corresponding restricted orbitals having maximal weight in the pairs of neighbor (overlapping) $\psi_{m}^{\text{BS}}$s.